Circular No. 7221

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990712 J. Heise and J. in ’t Zand, Space Research Organization of the Nether-

lands, Utrecht; G. Tarei and V. Torroni, BeppoSAX Science Operation Center, Telespazio, Rome; M. Feroci and G. Gandolfi, Istituto di Astrofisica Spaziale, CNR, Rome; and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, Bologna, report: “A γ-ray burst of moderate intensity, with the strongest x-ray counterpart observed to date, triggered the BeppoSAX Gamma-Ray Burst Monitor on July 12.69655 UT and was simultaneously detected in the Wide Field Camera (WFC) unit 2. The burst lasted for ∼ 30 s in both the γ-ray (40–700 keV) and x-ray (2–26 keV) energy ranges and has a double-peaked structure. The WFC-determined position of the x-ray counterpart to GRB 990712 is α = 22^{h}31^{m}50^{s}, δ = −73^{o}24′.4 (equinox 2000.0), with an error radius of 2′ (99-percent confidence level). Due to sun constraints, BeppoSAX cannot perform a follow-up observation with the Narrow Field Instruments.”

COMET C/1998 T1 (LINEAR)

Ephemeris from the orbital elements on MPEC 1999-M35:

1999 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
July 1 19^{h}57.^{m}53 −44^{o}38.^{′}0 0.490 1.470 153.^{o}1 18.^{o}2  8.1
11 16 12.50 −43 41.3 0.574 1.485 136.1 28.4  8.5
21 14 39.72 −32 49.5 0.836 1.512 109.1 39.4  9.4
31 14 07.33 −26 26.9 1.151 1.552  91.3 40.8 10.2
Aug. 10 13 53.87 −22 56.9 1.474 1.603  77.8 38.2 10.9
20 13 48.23 −20 55.9 1.787 1.663  66.4 33.9 11.5
30 13 46.48 −19 44.2 2.081 1.732  56.0 28.9 12.0
Sept. 9 13 46.93 −19 02.3 2.351 1.807  46.2 23.7 12.4
19 13 48.73 −18 39.5 2.594 1.888  36.8 18.6 12.8

Total visual magnitude estimates: July 8.52 UT, 9.2 (M. Mattiazzo, Wallaroo, South Australia, 0.20-m reflector); 9.39, 8.4 (S. T. Rae, Palmerston North, New Zealand, 0.11-m reflector); 10.31, 9.5 (N. Biver, Oahu, HI, 0.26-m reflector); 11.90, 8.6 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars); 13.54, 8.5 (Rae, 10×50 binoculars).

1999 July 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green